My answer(s) were marked as incorrect. My textbook's Ka values for H3PO4 are: Ka1=7.11x10^-3, Ka2=6.32x10^-8, and Ka3=4.5x10^-13What is the pH of the buffer formed by mixing 83.0 mL of 0.280 M Na H2PO4 with (This problem requires values in your textbook
Answer #1

H₂ POL kag = 4.5x1013 kan = 711x103 pice = -log(711x103) pka, = 2.148 - ka,- 6.32 X108 pka, z-609 (6-32x170) pka2 = 7.2 Pragainiticul 6 44 me of 0.47 m NaOH = NNU OH = (uu xo 47) = 20.68 mmol NAH2 POU + NaOH - Naz H POY + H2O (mmol) 23.24 20.68 2.56